# On the Sylvester program and Cayley algorithm for vector partition reduction

### v2

@@ 5 Vector partition with collinear columns
The example of partitions with collinear columns is discussed in Appendix C.
6 Computational aspects of partition reduction
In Section 4 we consider procedures required for evaluation of a contribution $C_{\bf s}^{1}$ to a sum of $m$ terms terms $W^{l}_{i}$ (50) or $\bar{W}^{l}_{i}$ (56) representing a reduction of $l$-tuple partition $W({\bf s},{\bf D})$ for nonnegative $l\times m$ matrix ${\bf D}$
$W({\bf s},{\bf D})=\sum_{i=1}^{m}W^{l}_{i}=\sum_{i=1}^{m}\bar{W}^{l}_{i}$
When the columns of ${\bf D}$ are noncollinear the contributions $C_{\bf s}^{i}$ are independent and $W({\bf s},{\bf D})$ can aslo be presented as a mixture of $W^{l}_{i}$ and $\bar{W}^{l}_{i}$.
In Section 4 the contribution $C_{\bf s}^{1}$ emerging due to elimination of the last row and the first column ${\bf c}_{1}$ of the augmented matrix ${\bf E}$ requires computation of the numerical matrix ${\bf D}_{1}$ and the vectorial arguments ${\bf s}_{1}$ of $W^{l}_{1}$ in (50) and ${\bf s}_{1}({\bf j})$ of $\bar{W}^{l}_{1}$ in (56). First focus on the matrix ${\bf D}_{1}$ and note that the $k$-th element of ${\bf c}^{1}_{j}$ column is a $2\times 2$ minor determinant ${\cal D}^{lk}_{1j}$ of a matrix ${\bf D}^{lk}_{1j}$ with the elements that belong to the $k$-th and $l$-th rows and columns ${\bf c}_{j}$ and ${\bf c}_{1}$ of the matrix ${\bf E}$. The zeroth column ${\bf c}_{0}$ is employed in the computation of ${\bf s}_{1}$ found as ${\bf s}_{1}={\cal D}^{lk}_{10}$. Computational procedures for $W^{l}_{i}$ and $\bar{W}^{l}_{i}$ for $i$ in the range $1\leq i\leq m$ are described below.
6.1 Compact reduction
Introduce an operation $\mathscr{R}_{i}^{lk}$ acting on $l\times(m+1)$ augmented matrix ${\bf E}$ consisting of the following steps
  1. 1.
[…]
    ${\bf E}_{i}=\{\bm{r}_{1i},\bm{r}_{2i},\ldots,\bm{r}_{li}\}.$
  2. 2.
    Select $k$-th and $l$-th rows $\bm{r}_{ki},\bm{r}_{li}$ with $k<l$ of ${\bf E}_{i}$ and form $2\times 2$ matrices from the elements belonging to these two rows and column ${\bf c}_{j}$ and $k$-th and $l$-th elements $c_{ik},c_{il}$ of column ${\bf c}_{i}$
    ${\bf D}^{lk}_{ij}=\left(\begin{array}[]{cc}c_{jk}&c_{ik}\\ c_{jl}&c_{il}\end{array}\right),\quad 0\leq j\leq m,\ j\neq i.$
  3. 3.
    Compute the minor determinants ${\cal D}^{lk}_{ij}=\det{\bf D}^{lk}_{ij}$.
  4. 4.
    Form a row ${\cal R}_{i}^{k}$ with $m$ elements ${\cal D}^{lk}_{ij}$
    ${\cal R}_{i}^{k}=\mathscr{R}_{i}^{lk}[{\bf E}]=\{{\cal D}^{lk}_{i0},{\cal D}^{lk}_{i1},\ldots,{\cal D}^{lk}_{i,i-1},{\cal D}^{lk}_{i,i+1},\ldots,{\cal D}^{lk}_{im}\}.$
Now define an operation $\mathscr{R}_{i}^{l}[{\bf E}]$ that produces an array of ${\cal R}_{i}^{k}$ for $11\leq k\leq l-$ 1
$\mathscr{R}_{i}^{l}[{\bf E}]=\{{\cal R}_{i}^{1},{\cal R}_{i}^{2},\ldots,{\cal R}_{i}^{l-1}\},$
that represents a new matrix ${\cal E}_{i}=\mathscr{R}_{i}^{l}[{\bf E}]$. After application of the transformations (51) and (52) to ${\cal E}_{i}$ we obtain a new matrix ~$\tilde{\cal E}_{i}$ with the last row free of negative elements. The term $W_{i}^{l}({\bf s})$ is given by
$W_{i}^{l}({\bf s})=(-1)^{K_{i}}H(\tilde{s}^{i}_{l-1})W(\tilde{\cal E}_{i}),$ (74)
where $K_{i}$ is the number of negative elements in the last row of the matrix ${\cal E}_{i}$ and $\tilde{s}^{i}_{l-1}$ denotes the last element of the argument column $\tilde{\bf s}^{i}$ of the matrix ~$\tilde{\cal E}_{i}$.
6.2 Expanded reduction
Consider the case of a matrix ${\bf E}$ containing NP-columns with GCD larger than unity $\mbox{gcd}({\bf c})_{i}>1$ when the contribution of such column ${\bf c}_{i}$ given by a superposition $\bar{W}^{l}_{i}({\bf s})$ of SPFs.
The only difference to the compact reduction case is the zeroth column
${\bf s}_{i}({\bf j})=({\bf s}_{i}-\hat{\bf j}c_{il}+j_{l}\hat{\bf c}_{i})/c_{il},\quad{\bf s}_{i}=\{{\cal D}^{l1}_{i0},{\cal D}^{l2}_{i0},\ldots,{\cal D}^{l,l-1}_{i0}\}=c_{il}\hat{\bf s}-s_{l}\hat{\bf c}_{i}.$
We observe that the expression $c_{il}{\bf s}_{i}({\bf j})=c_{il}(\hat{\bf s}-\hat{\bf j})-(s_{l}-j_{l})\hat{\bf c}_{i}$ coincides with ${\bf s}_{i}$ in (56) where ${\bf s}$ is replaced by ${\bf s}-{\bf j}$, and thus the expression for $\bar{W}_{i}^{l}({\bf s})$ reads
$\bar{W}_{i}^{l}({\bf s})=\!\!\!\!\sum_{{\bf j}={\bf N}_{i}^{-}}^{{\bf N}_{i}^{+}}a_{i,{\bf j}}W({\cal E}_{i}({\bf j}))\delta_{j_{l},s_{l}\bmod{c_{il}}},\quad{\cal E}_{i}({\bf j})=\mathscr{R}_{i}^{l}[{\bf E}_{i}({\bf j})],\quad{\bf E}_{i}({\bf j})=\{({\bf s}-{\bf j})/c_{il},{\bf c}_{1},{\bf c}_{2},\ldots,{\bf c}_{m}\}.$ (75)
Comparing (75) to (74) we observe that $\bar{W}_{i}^{l}$ can be viewed as a weighted sum of $W_{i}^{l}$ with shifted argument
$\bar{W}_{i}^{l}({\bf s})=\!\!\!\!\sum_{{\bf j}={\bf N}_{i}^{-}}^{{\bf N}_{i}^{+}}a_{i,{\bf j}}W_{i}^{l}(({\bf s}-{\bf j})/c_{il})\delta_{j_{l},s_{l}\bmod{c_{il}}}.\quad\sum_{{\bf j}={\bf N}_{i}^{-}}^{{\bf N}_{i}^{+}}a_{i,{\bf j}}=c_{il}^{m-1},$ (76)
where the summation limits are determined by the following procedure. Introduce reduced rows $\bm{r}^{\prime}_{ji}$ each having $(m-1)$ elements by dropping first element of $\bm{r}_{ji}$ and define their $L_{1}$-norm $|\bm{r}^{\prime}_{ji}|=\sum_{k=1}^{m}c_{kj},\ k\neq i$. For each row $\bm{r}^{\prime}_{ji}$ define $(l-1)$ quantities
$N_{ji}^{-}=\min(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{ji}-c_{ij}\lfloor(\bm{K}^{\prime}_{p}\cdot\bm{r}^{\prime}_{li})/c_{il}\rfloor),\quad N_{ji}^{+}=|\bm{r}^{\prime}_{ji}|,\quad 1\leq j\leq l-1,$ (77)
where
$\bm{K}^{\prime}_{p}=\{k_{1},k_{2},\ldots,k_{i-1},k_{i+1},\ldots,k_{m}\},\quad 0\leq k_{i}\leq c_{il}-1,\quad 1\leq p\leq c_{il}^{m-1},$
and form two vectors ${\bf N}_{i}^{\pm}$ representing the summation limits in (76)
${\bf N}_{i}^{-}=\{N_{1i}^{-},N_{2i}^{-},\ldots,N_{l-1,i}^{-},0\}^{T},\quad{\bf N}_{i}^{+}=(c_{il}-1)\{N_{1i}^{+},N_{2i}^{+},\ldots,N_{l-1,i}^{+},1\}^{T}.$ (78)
To find the lower bound of $N_{ji}^{-}$ note that $N_{ji}^{-}\geq\bm{K}^{\prime}_{p}\cdot(\bm{r}^{\prime}_{ji}-\bm{r}^{\prime}_{li})=\bm{K}^{\prime}_{p}\cdot\delta\bm{r}^{\prime}_{jl,i}.$ Replace all positive elements of the vector $\delta\bm{r}^{\prime}_{jl,i}$ by zeros producing a vector $\widehat{\delta\bm{r}}^{\prime}_{jl,i}$ and compute its $L_{1}$-norm $|\widehat{\delta\bm{r}}^{\prime}_{jl,i}|$. Then
$N_{ji}^{-}\geq(c_{il}-1)|\widehat{\delta\bm{r}}^{\prime}_{jl,i}|.$
6.3 Reduction of triple partition always leads to NP-columns
Consider triple partition transformation to SPFs that requires two reduction steps. Assuming that the corresponding matrix ${\bf E}$ is free of NP-columns we show below that among all two-row matrices generated by the reduction algorithm there should be a matrix with at least one NP-column.
Perform elimination of the last third row and first column ${\bf c}_{1}=\{c_{11},c_{12},c_{13}\}^{T}$ that produces a two-row matrix ${\cal E}_{1}$ with at least three numerical columns. Consider three columns ${\bf c}^{1}_{p}=\{{\cal D}^{31}_{1p},{\cal D}^{32}_{1p}\}^{T},\ {\bf c}^{1}_{r}=\{{\cal D}^{31}_{1r},{\cal D}^{32}_{1r}\}^{T},$ and ${\bf c}^{1}_{s}=\{{\cal D}^{31}_{1s},{\cal D}^{32}_{1s}\}^{T}$ of this matrix that computed using three pairs of columns $\{{\bf c}_{j},{\bf c}_{1}\},\ j=p,r,s$. We are interested in ${\bf c}^{1}_{j}$ with at least one odd element otherwise they are NP-columns automatically.
Below we review several cases of column elements parity and from this point to the end of this subsection all column elements are written modulo 2.
  1. 1.
    Assume that all ${\bf c}^{1}_{j}$ have both odd elements.
    1. (a)
      $c_{13}=1$ and $c_{j3}=1$. We find for $c_{11}=x$ and $c_{12}=y$ we have to have $c_{j1}=1-x$ and $c_{j1}=1-y$. Then elimination of any of ${\bf c}_{j}$ leads to a NP-column as, for example, ${\bf c}^{r}_{p}=\{{\cal D}^{31}_{rp},{\cal D}^{32}_{rp}\}^{T}$ has only even elements.
    2. (b)
      $c_{13}=1$ and $c_{j3}=0$. We find $c_{j1}=c_{j2}=1$ and again elimination of any of ${\bf c}_{j}$ leads to a NP-column.
    3. (c)
      $c_{11}=c_{12}=1,\ c_{13}=0$ and $c_{j3}=1$ that leads to $c_{p1}=c_{p2}=x$ and $c_{r1}=c_{r2}=1-x$ and ${\bf c}_{s}$ is equal to either ${\bf c}_{r}$ or ${\bf c}_{p}$, and elimination of ${\bf c}_{s}$ leads to a NP-column.
  2. 2.
    Assume that ${\bf c}^{1}_{p}$ and ${\bf c}^{1}_{r}$ have both odd elements while ${\bf c}^{1}_{s}$ has at least one even element. We have to consider only the case $c_{11}=c_{12}=1,\ c_{13}=0$, $c_{p3}=c_{r3}=1$ and $c_{p1}=c_{p2}=x$ with $c_{r1}=c_{r2}=1-x$ leading to ${\bf c}^{r}_{p}$ and ${\bf c}^{p}_{r}$ having both odd elements. In order to have one even element in ${\bf c}^{1}_{s}$ we demand $c_{s3}=0$ so that the second element is also even and ${\bf c}^{1}_{s}$ ends up to be a NP-column.
It is clear that reduction of vector partition defined by a $l$-row augmented matrix ${\bf E}$ with $l>2$ always produces NP-columns that must be processed using (76). In Appendix A we discuss the reduction of a triple partition into a sum of double partitions all them having a NP-column.
In Section 4.2.2 we mention that elimination of $i$-th NP-column produces matrices ${\bf D}_{i}$ with the numerical elements having the same common factor $g_{i}$ and one can remove this factor by dividing the whole matrix ${\bf D}_{i}$ by $g_{i}$. In some cases the matrix of three or more rows does not have NP-columns but its elements are such that the elimination of some columns leads to matrices where not the whole matrix but some rows have GCD larger than unity. In this case it is convenient to simplify each such row by dividing it on GCD of its numerical elements.
Using the row simplification method, we find that the reduction of a triple partition in some rare cases eventually leads to two-row matrices void of NP-columns; the reduction of a triple partition in Appendix B gives such an example.
6.4 Iterative reduction of VPF to SPFs
@@
The basic step of iterative reduction procedure of the vector partition corresponding to the $l$-row matrix ${\bf E}$ described in Section 4 and above in this Section includes elimination of the last row and a single $i$-th column of ${\bf E}$. It results in a single partition based on the smaller matrix ~$\tilde{\cal E}_{i}$ for P-column or a superposition of VPFs built on matrices ~$\tilde{\cal E}_{i}({\bf j})$ for NP-column. Each such matrix undergoes similar transformation in the consecutive iterative steps leading to combination of scalar partitions. The total number of iterations required to reach this stage is $l-1$.
@@
We note in Section 3.1 that the order of the rows in the two-row matrix ${\bf E}$ affects the individual scalar partition contributions to the double partition. The same holds for a matrix with the larger number of rows. This means that while one can change row order in the initial matrix ${\bf E}$, it is forbidden for any matrix generated by the iterative elimination procedure.
For definiteness assume that the order of row elimination is fixed and at every step we eliminate the last row of each generated matrix. For some matrices ${\bf E}$ this elimination order leads to appearance of zero elements in the last row of the generated matrices of smaller dimensions. This case requires a special treatment that discussed in Section 4.3.
@@
How can one define the partition function in this case? The last row ${\bf R}_{n}$ of the matrix corresponds to a scalar partition defined by (1) that requires to find all sets ${\bf x}_{K}$ of nonnegative integer solutions. Assume that we find all such sets, their number is given by a scalar partition $W(S_{n},\{R_{n1},\ldots,R_{np}\})$. Now substitute each solution ${\bf x}_{K}$ into the remaining $n-1$ equations and retain only those that satisfy them all. The number of such solutions is the vector partition $W({\bf E}_{\nu})$. This definition fails when the last row has zero entries as the number of solutions ${\bf x}_{K}$ in this case is infinite. Nevertheless, one still can proceed with the evaluation procedure in these cases, this problem is addressed in Section 4.3 where we prove that the contribution of a column with last zero element is zero.
In the original definition of the vector partition mentioned in Section 1 the elements of the vector ${\bf s}$ are considered nonnegative. This assumption does not hold in general for ${\bf E}_{\nu}$ as the linear superpositions $L_{k}({\bf s})$ might be negative for some values of the vector ${\bf s}$. For the rows with negative numerical elements negative $L_{k}({\bf s})<0$ are acceptable. But if a row ${\bf R}_{k}$ is void of negative generators we have to restrict its element $L_{k}({\bf s})$ to nonnegative values $L_{k}({\bf s})\geq 0$. This requirement forces us to replace the partition $A_{\nu}W({\bf E}_{\nu})$ by
$A_{\nu}W({\bf E}_{\nu})\Rightarrow A_{\nu}B_{\nu}W({\bf E}_{\nu}),\quad B_{\nu}=\prod_{k}H(L_{k}({\bf s})).$ (79)
@@ 7 Conclusive remarks
The results presented in the manuscript show that Sylvester strongly underestimated the number of scalar partitions required for computation of a vector partition – their minimal number in case of $l\times m$ matrix significantly surpasses $m!/(m-l+1)!$ which in its turn is $(l-1)!$ times larger than predicted by Sylvester in [12]. We address the case of collinear columns of the generator matrix in Section 5 and show that in this case a vector partition boils down to a convolution of scalar partitions.
Our approach is a generalization of the algorithms discussed in [9] and we demonstrate that a vector partition always reduces iteratively either to a sum of scalar partitions themselves or their convolutions, and thus the Sylvester program is successfully implemented.
Recalling that the scalar partitions can be computed as a superposition of the known polynomials [6] we arrive at the important statement — any vector partition can be expressed through the Bernoulli polynomials of higher order.
@@ Appendix A Triple partition reduction example
where each matrix ${\bf E}_{i}$ contains the NP-column. Perform the following transformation of the matrices ${\bf E}_{i}$ – first divide each row by the GCD of its numerical elements, and then use (51, 52) to make last row positive, and obtain a sum of four double partitions
@@
As discussed in Section 6.4 each summand in (A13) acquires an additional factor producing
@@ Appendix B Alternative triple partition reduction
$W({\bf E}^{\prime})=\sum_{k=1}^{4}A^{\prime}_{k}W(\tilde{\bf E}^{\prime}_{k}),\quad$ (B13)
where
$A^{\prime}_{1}=-H(s_{1}-s_{3}-5),\quad A^{\prime}_{2}=-A^{\prime}_{3}=H(3s_{1}-s_{3}),\quad A^{\prime}_{4}=H(2s_{1}-s_{3}-2).$ (B14)
@@
The total number of scalar partitions in (B15) is ten which is significantly smaller than we have for the triple partition considered in Appendix A.
@@ Appendix C Partition with collinear columns
Consider reduction of a triple partition $W({\bf E})$ defined by the matrix
@@
that replaces $W({\bf S}_{3}({\bf k}),\bar{\bf D}_{3})$ in (C29) to find explicit expression for $W({\bf s}_{4},{\bf D}_{4})$. Finally, consider $W({\bf s}_{5},{\bf D}_{5})$ with the matrix ${\bf D}_{5}$ void of noncollinear columns. This case is served by the relations (72,73) derived in Section 5. We choose the vectors ${\bf C}_{p}$ as follows ${\bf C}_{1}=\{2,1\}^{T}$ with ${\bf u}_{1}=\{1,1\}$; ${\bf C}_{2}=\{1,0\}^{T}$ with ${\bf u}_{2}=\{1,1\}$; ${\bf C}_{3}=\{0,1\}^{T}$ with ${\bf u}_{3}=\{1,1\}$, and ${\bf C}_{4}=\{1,1\}^{T}$ with ${\bf u}_{4}=\{1,1,1,1\}$. Use the vectors ${\bf s}_{5}=\{s_{3}-s_{1}-1,s_{2}-s_{1}-2\},\ {\bf s}_{6}=\{s_{3}-2s_{1}-4,s_{2}-s_{1}-2\}$ to find

### v3

@@ 5 Vector partition with collinear columns
The example of partitions with collinear columns is discussed in Appendix C where we describe the reduction procedure for the matrix used in computation of the Kronecker coefficients [9, 10].
6 Iterative reduction of vector to scalar partitions
@@
The basic step of iterative reduction procedure of the vector partition corresponding to the $l$-row matrix ${\bf E}$ described in Section 4 includes elimination of the last row and a single $i$-th column of ${\bf E}$. It results in a single partition based on the $(l-1)$-row matrix for the P-column or a superposition of partitions built on similar matrices for the NP-column. Each such matrix undergoes similar transformation in the consecutive iterative steps leading to a combination of scalar partitions. The total number of iterations required to reach this stage is $l-1$.
@@
We note in Section 3.1 that the order of the rows in the two-row matrix ${\bf E}$ affects the individual scalar partition contributions to the double partition. The same holds for a matrix with the larger number of rows. This means that while one can change row order in the initial matrix ${\bf E}$, it is forbidden for any matrix generated in the course of the iterative elimination procedure. For definiteness assume that the order of row elimination is fixed and at every step we eliminate the last row of each generated matrix.
@@
How can one define the partition function in this case? The last row ${\bf R}_{n}$ of the matrix corresponds to a scalar partition defined by (1) that requires to find all sets ${\bf x}_{K}$ of nonnegative integer solutions. Assume that we find all such sets, their number is given by a scalar partition $W(S_{n},\{R_{n1},\ldots,R_{np}\})$. Now substitute each solution ${\bf x}_{K}$ into the remaining $n-1$ equations and retain only those that satisfy them all. The number of such solutions is the vector partition $W({\bf E}_{\nu})$. This definition fails when the last row has zero entries as the number of solutions ${\bf x}_{K}$ in this case is infinite. Nevertheless, one still can proceed with the evaluation procedure in these cases, this problem is addressed in Section 4.3 where we prove that the contribution of a column with the last zero element is zero.
In the original definition of the vector partition mentioned in Section 1 the elements of the vector ${\bf s}$ are considered nonnegative. This assumption does not hold in general for ${\bf E}_{\nu}$ as the linear superpositions $L^{\prime}_{k}({\bf s})$ might be negative for some values of the vector ${\bf s}$. For the rows with negative numerical elements the negative $L^{\prime}_{k}({\bf s})<0$ are acceptable. But if the row ${\bf R}_{k}$ is void of negative numbers we have to restrict its initial element $L^{\prime}_{k}({\bf s})$ to the nonnegative values $L^{\prime}_{k}({\bf s})\geq 0$ and to replace the partition $A_{\nu}W({\bf E}_{\nu})$ by
$A_{\nu}W({\bf E}_{\nu})\Rightarrow A_{\nu}B_{\nu}W({\bf E}_{\nu}),\quad B_{\nu}=\prod_{k}H(L^{\prime}_{k}({\bf s})).$ (70)
@@ 7 Conclusive remarks
The results presented in the manuscript show that Sylvester strongly underestimated the number of scalar partitions required for computation of the vector partition – their minimal number in case of $l\times m$ positive matrix might significantly surpass $m!/(m-l+1)!$ which in its turn is $(l-1)!$ times larger than predicted by Sylvester in [17]. We address the case of collinear columns of the generator matrix in Section 5 and show that in this case a vector partition boils down to a convolution of scalar partitions.
Our approach is a generalization of the algorithms discussed in [14] and here we demonstrate that a vector partition always reduces iteratively either to a sum of scalar partitions themselves or their convolutions, and thus the Sylvester program is successfully implemented. Recalling that the scalar partitions can be computed as a superposition of the known polynomials [11] we arrive at the important statement — any vector partition can be expressed through the Bernoulli polynomials of higher order.
All known approaches to evaluation of vector partitions (see [1, 15, 3, 18]), their later modifications [2, 5, 8] and the software packages based of these methods successfully determine the chambers but fail to produce the explicit expressions for the vector partitions as they replace the original generating function by the signed sum of the simpler partial generating functions valid in specific chamber. This means that for each new argument column ${\bf s}$ one has to perform the Taylor expansion of the partial generating functions to extract the desired coefficient. Our method is the first one that permits the derivation of a single explicit formula for the vector partition which is valid for an arbitrary value of ${\bf s}$. The usage of this formula does not require the explicit knowledge of the chambers but nevertheless it implicitly contains their description.
@@ Appendix A Triple partition reduction example
where each matrix ${\bf E}_{i}$ contains the NP-column. Perform the following transformation of the matrices ${\bf E}_{i}$ – first divide each row by the GCD of its numerical elements, and then use (46, 47) to make last row positive, and obtain a sum of four double partitions
@@
As discussed in Section 6 each summand in (A13) acquires an additional factor producing
@@ Appendix B Alternative triple partition reduction
$W({\bf E}^{\prime})=\sum_{k=1}^{4}A^{\prime}_{k}W(\tilde{\bf E}^{\prime}_{k}),\ A^{\prime}_{1}=-H(s_{1}-s_{3}-5),\ A^{\prime}_{2}=-A^{\prime}_{3}=H(3s_{1}-s_{3}),\ A^{\prime}_{4}=H(2s_{1}-s_{3}-2).$
@@
The total number of scalar partitions in (B13) is ten which is significantly smaller than we have in Appendix A.
@@ Appendix C Partition with collinear columns
Consider reduction of a triple partition $W({\bf E})$ defined by the matrix used in the computation of the Kronecker coefficients [9, 10]
@@
that replaces $W({\bf S}_{3}({\bf k}),\bar{\bf D}_{3})$ in (C28) to find explicit expression for $W({\bf s}_{4},{\bf D}_{4})$. Finally, consider $W({\bf s}_{5},{\bf D}_{5})$ with the matrix ${\bf D}_{5}$ void of noncollinear columns. This case is served by the relations (68,69) derived in Section 5. We choose the vectors ${\bf C}_{p}$ as follows ${\bf C}_{1}=\{2,1\}^{T}$ with ${\bf u}_{1}=\{1,1\}$; ${\bf C}_{2}=\{1,0\}^{T}$ with ${\bf u}_{2}=\{1,1\}$; ${\bf C}_{3}=\{0,1\}^{T}$ with ${\bf u}_{3}=\{1,1\}$, and ${\bf C}_{4}=\{1,1\}^{T}$ with ${\bf u}_{4}=\{1,1,1,1\}$. Use the vectors ${\bf s}_{5}=\{s_{3}-s_{1}-1,s_{2}-s_{1}-2\},\ {\bf s}_{6}=\{s_{3}-2s_{1}-4,s_{2}-s_{1}-2\}$ to find